# Linear Bandits with Partially Observable Features

### v1

@@ C.2 Proof of Theorem 2
$\left\|\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\left(\check{\mathbf{w}}_{t}^{L}-\mathbf{w}_{t}\right)\right\|_{\infty}\leq\frac{4\sigma\epsilon}{1-\epsilon}\sqrt{\frac{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}}{p}}.$
Now we obtain,
$\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}(\widetilde{y}_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\mathbf{w}_{t})\mathbf{V}_{t}^{-1/2}\widetilde{\mathbf{x}}_{a}\right\|_{\infty}$ $\leq\frac{4\sigma\epsilon}{1-\epsilon}\sqrt{\frac{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}}{p}}+\frac{\sigma}{p}\sqrt{2\log\frac{2Kt^{2}}{\delta}}$
$\leq\frac{4\sigma\epsilon}{1-\epsilon}\sqrt{\frac{2K\log\frac{2Kt^{2}}{\delta}}{p}}+\frac{\sigma}{p}\sqrt{2\log\frac{2Kt^{2}}{\delta}}$
$=\left(\frac{4\epsilon\sqrt{K}}{1-\epsilon}+\frac{1}{p}\right)\sigma\sqrt{2\log\frac{2Kt^{2}}{\delta}}$
Setting $\epsilon=p^{-1}/(\sqrt{K}+p^{-1})$ gives $\frac{\epsilon\sqrt{K}}{{1-\epsilon}}=p^{-1}$ and for $t\geq 8(\sqrt{K}+p^{-1})^{2}p^{2}(1-p)^{-2}K^{2}\log\frac{2Kt^{2}}{\delta}$,
$\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}(\widetilde{y}_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\mathbf{w}_{t})\mathbf{V}_{t}^{-1/2}\widetilde{\mathbf{x}}_{a}\right\|_{\infty}\leq\frac{5\sigma}{p}\sqrt{2\log\frac{2Kt^{2}}{\delta}}$
Because $\widehat{\mathbf{w}}_{t}$ is a minimizer of (11), by Lemma 3,
$\left\|\widehat{\mathbf{w}}_{t}-\mathbf{w}_{t}\right\|_{\mathbf{V}_{t}^{-1/2}(\sum_{\tau=1}^{t}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a})\mathbf{V}_{t}^{-1/2}}\leq\frac{20\sigma}{p}\sqrt{\frac{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}}{\lambda_{\min}\left(\mathbf{V}_{t}^{-1/2}(\sum_{\tau=1}^{t}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a})\mathbf{V}_{t}^{-1/2}\right)}},$
which is equivalent to,
$\left\|\widehat{\mathbf{w}}_{t}-\mathbf{w}_{t}\right\|_{2}\leq\frac{20\sigma}{p}\sqrt{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}}.$
This concludes,
$\max_{a\in[K]}|\widetilde{\mathbf{x}}_{a}(\widehat{\boldsymbol{\mu}}_{t}-\boldsymbol{\mu}_{\star})|\leq\frac{20\sigma}{p}\sqrt{\frac{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}}{t}},$
which conmpletes the proof. ∎
C.3 Proof of Theorem 3
Because the regret is bounded by $2$ and the number of rounds for the exploration phase is at most $|\mathcal{E}_{T}|\leq 8(\sqrt{K}+p^{-1})^{2}p^{2}(1-p)^{-2}K^{2}\log\frac{2KT^{2}}{\delta}$.
$\mathrm{Reg}(T)$ $\leq\frac{16K^{2}(\sqrt{K}+p^{-1})^{2}}{(1-p)^{2}}\log\frac{2Kt^{2}}{\delta}+\sum_{t\in[T]\setminus\mathcal{E}_{T}}\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]$
$=$ $\frac{16K^{2}(\sqrt{K}+p^{-1})^{2}}{(1-p)^{2}}+\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}\left(a_{t}=\widehat{a}_{t}\right)\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}$
@@ C.3 Proof of Theorem 3
$\leq\frac{40\sigma}{p}\sqrt{\frac{2(d+d_{h})}{t}\log\frac{2Kt^{2}}{\delta}},$
@@
$\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}\left(a_{t}=\widehat{a}_{t}\right)\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}\leq\frac{80\sigma}{p}\sqrt{\frac{2(d+d_{h})}{t}\log\frac{2Kt^{2}}{\delta}}.$
By resampling at most $\rho_{t}$ times, the probability of the event $\{a_{t}\neq\widehat{a}_{t}\}$ is
$\mathbb{P}\left(a_{t}\neq\widehat{a}_{t}\right)=$ $\sum_{m=1}^{\rho_{t}}\frac{p}{(K-1)\sqrt{t}}\left(1-\frac{p}{(K-1)\sqrt{t}}\right)^{m-1}\cdot\mathbb{P}(\text{Resample succeed at trial }m)$
$\leq$ $\sum_{m=1}^{\rho_{t}}\frac{p}{(K-1)\sqrt{t}}\left(1-\frac{p}{(K-1)\sqrt{t}}\right)^{m-1}$
$=$ $\frac{p}{(K-1)\sqrt{t}}\left(\frac{p}{(K-1)\sqrt{t}}\right)^{-1}\left\{1-\left(1-\frac{p}{(K-1)\sqrt{t}}\right)^{\rho_{t}}\right\}$
$=$ $1-\left(1-\frac{p}{(K-1)\sqrt{t}}\right)^{\rho_{t}}$
$\leq$ $\frac{p\rho_{t}}{(K-1)\sqrt{t}},$
where the last inequality uses $(1+x)^{n}\geq 1+nx$ for $x\geq-1$ and $n\in\mathbb{N}$. Then the expected sum of regret,
$\mathbb{E}\left[\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}\left(a_{t}=\widehat{a}_{t}\right)\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}\right]$
$\leq 2\sum_{t\in[T]\setminus\mathcal{E}_{T}}\mathbb{P}\left(a_{t}\neq\widehat{a}_{t}\right)$
$\leq\frac{4p\sqrt{T}}{K-1}\rho_{T}$
$=\frac{4p\sqrt{T}}{K-1}\frac{\log\frac{(T+1)^{2}}{\delta}}{\log(1/p)}.$
Thus,
$\mathbb{E}[\mathrm{Reg}(T)]$ $\leq 10\delta\log T+\frac{16K^{2}(\sqrt{K}+p^{-1})^{2}}{(1-p)^{2}}+\frac{4p\sqrt{T}}{K-1}\frac{\log\frac{(T+1)^{2}}{\delta}}{\log(1/p)}$
$\quad+\frac{80\sigma}{p}\sqrt{2(d+d_{h})T\log\frac{2KT^{2}}{\delta}},$
@@ C.4 Proof of Theorem 4
$\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}-{\boldsymbol{\mu}}_{\star})=\widetilde{\mathbf{x}}_{a}^{\top}\widetilde{\mathbf{V}}_{t}^{-1}\left\{\sum_{\tau=1}^{t}\mathbb{I}(\mathcal{M}_{\tau})\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\left(\widetilde{y}_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}{\boldsymbol{\mu}}_{\star}\right)-{\boldsymbol{\mu}}_{\star}\right\}.$
@@
$\widetilde{y}_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}{\boldsymbol{\mu}}_{\star}=\left(1-\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,t}}\right)\widetilde{\mathbf{x}}_{a}^{\top}\left(\check{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star}\right)+\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\epsilon_{a,\tau}.$
@@
$\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}-{\boldsymbol{\mu}}_{\star})=\widetilde{\mathbf{x}}_{a}^{\top}\widetilde{\mathbf{V}}_{t}^{-1}\left\{\left(\widetilde{\mathbf{V}}_{t}-\widetilde{\mathbf{A}}_{t}\right)\left(\check{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star}\right)+\sum_{\tau=1}^{t}\mathbb{I}(\mathcal{M}_{\tau})\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\widetilde{\mathbf{x}}_{a}\epsilon_{a,\tau}-{\boldsymbol{\mu}}_{\star}\right\}.$
@@
$\check{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star}=$ $\left(\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}+p\mathbf{I}_{K}\right)^{-1}\left(\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{a_{\tau},\tau}-p{\boldsymbol{\mu}}_{\star}\right)$
$=$ $\left(\sum_{\tau=1}^{t}\frac{1}{\phi_{a_{\tau},\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K}\right)^{-1}\left(\sum_{\tau=1}^{t}\frac{1}{\phi_{a_{\tau},\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{a_{\tau},\tau}-{\boldsymbol{\mu}}_{\star}\right)$
$=\left(\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K}\right)^{-1}\left(\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{a_{\tau},\tau}-{\boldsymbol{\mu}}_{\star}\right),$
@@
$\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}-{\boldsymbol{\mu}}_{\star})=$ $\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1}\left\{\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{A}_{t}^{-1}\left(\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{a_{\tau},\tau}-{\boldsymbol{\mu}}_{\star}\right)+\sum_{\tau=1}^{t}\frac{\epsilon_{a_{\tau}}}{\phi_{a_{\tau},\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}-{\boldsymbol{\mu}}_{\star}\right\}$
$=$ $\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1}\left\{\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{A}_{t}^{-1}+\mathbf{I}_{K}\right\}\left(\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{a_{\tau},\tau}-{\boldsymbol{\mu}}_{\star}\right)$
$=$ $\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}^{1/2}\mathbf{A}_{t}^{-1}\mathbf{V}_{t}^{1/2}\right)\mathbf{V}_{t}^{-1/2}\left(\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{a_{\tau},\tau}-{\boldsymbol{\mu}}_{\star}\right).$
@@
$\max_{a\in[K]}|\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}-{\boldsymbol{\mu}}_{\star})|\leq\max_{a\in[K]}\|\widetilde{\mathbf{x}}_{a}\|_{\mathbf{V}_{t}^{-1}}\|\mathbf{V}_{t}^{1/2}\mathbf{A}_{t}^{-1}\mathbf{V}_{t}^{1/2}\|_{2}\left\|\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{a_{\tau},\tau}-{\boldsymbol{\mu}}_{\star}\right\|_{\mathbf{V}_{t}^{-1}}.$
@@
$\max_{a\in[K]}|\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}-{\boldsymbol{\mu}}_{\star})|$ $\leq\frac{\max_{a\in[K]}\|\widetilde{\mathbf{x}}_{a}\|_{\mathbf{V}_{t}^{-1}}}{1-\epsilon}\left\|\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{a_{\tau},\tau}-{\boldsymbol{\mu}}_{\star}\right\|_{\mathbf{V}_{t}^{-1}}$
$\leq\frac{\max_{a\in[K]}\|\widetilde{\mathbf{x}}_{a}\|_{\mathbf{V}_{t}^{-1}}}{1-\epsilon}\left(\frac{1}{p}\left\|\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{a_{\tau},\tau}\right\|_{\mathbf{V}_{t}^{-1}}+\left\|{\boldsymbol{\mu}}_{\star}\right\|_{\mathbf{V}_{t}^{-1}}\right).$
Note that the matrix $\mathbf{V}_{t}$ is deterministic. By Lemma 9 in [Abbasi-yadkori et al., 2011], with probability at least $1-\delta,$
$\left\|\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{a_{\tau},\tau}\right\|_{\mathbf{V}_{t}^{-1}}$ $\leq\left\|\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{a_{\tau},\tau}\right\|_{\left(\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K}\right)^{-1}}$
@@
$\left\|\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{a_{\tau},\tau}\right\|_{\mathbf{V}_{t}^{-1}}\leq\sigma\sqrt{K\log\frac{t+1}{\delta}},$
@@
$\max_{a\in[K]}|\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}-{\boldsymbol{\mu}}_{\star})|$ $\leq\frac{\max_{a\in[K]}\|\widetilde{\mathbf{x}}_{a}\|_{\mathbf{V}_{t}^{-1}}}{1-\epsilon}\left(\frac{\sigma}{p}\sqrt{K\log\frac{t+1}{\delta}}+\left\|{\boldsymbol{\mu}}_{\star}\right\|_{\mathbf{V}_{t}^{-1}}\right)$
@@
C.5 Proof of Theorem 5
Because the regret is bounded by $1$ and the number of rounds for the exploration phase is at most $|\mathcal{E}_{T}|\leq 32(1-p)^{-2}K^{2}\log\frac{2dT^{2}}{\delta}.$

### v2

@@ E.2 Proof of Theorem 2
$\left\|\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\left(\check{\boldsymbol{\mu}}_{t}^{L}-\boldsymbol{\mu}_{\star}\right)\right\|_{\infty}\leq\frac{4\epsilon\sigma\widetilde{\sigma}^{2}_{\max}}{p(1-\epsilon)\widetilde{\sigma}_{\min}}\sqrt{2t(d+d_{h})\log\frac{2Kt^{2}}{\delta}}.$
Setting $\epsilon=K^{-1/2}\widetilde{\sigma}_{\min}\widetilde{\sigma}_{\max}^{-1}/8$ gives $21-\epsilon\geq 1/$ 2 and
$\left\|\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\left(\check{\boldsymbol{\mu}}_{t}^{L}-\boldsymbol{\mu}_{\star}\right)\right\|_{\infty}$ $\leq\frac{\sigma\widetilde{\sigma}_{\max}}{p}\sqrt{2t\frac{d+d_{h}}{K}\log\frac{2Kt^{2}}{\delta}}$
$\leq\frac{\sigma\widetilde{\sigma}_{\max}}{p}\sqrt{2t\log\frac{2Kt^{2}}{\delta}}.$
Now we conclude that for $t\geq 8^{3}K^{3}\widetilde{\sigma}^{-2}_{\min}\widetilde{\sigma}_{\max}^{2}(1-p)^{-2}\log(2Kt^{2}/\delta)$
$\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}(\widetilde{y}_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}\boldsymbol{\mu}_{t})\widetilde{\mathbf{x}}_{a}\right\|_{\infty}\leq\frac{\sigma\widetilde{\sigma}_{\max}}{p}\sqrt{2t\log\frac{2Kt^{2}}{\delta}},$
and by Lemma 3,
$\|\widehat{\boldsymbol{\mu}}_{t}^{L}-\boldsymbol{\mu}_{\star}\|_{\mathbf{V}_{t}}\leq 8\sigma\widetilde{\sigma}_{\max}\sqrt{\frac{2t(d+d_{h})\log(2Kt^{2}/\delta)}{p\lambda_{\min}\left(\mathbf{V}_{t}\right)}}=\frac{8\sigma\widetilde{\sigma}_{\max}}{p\widetilde{\sigma}_{\min}}\sqrt{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}},$
which completes the proof. ∎
E.3 Proof of Theorem 3
Because the regret is bounded by $2$ and the number of rounds for the exploration phase is at most $|\mathcal{E}_{T}|\leq 8^{3}K^{3}(1-p)^{-2}\log(2KT^{2}/\delta)$.
$\mathrm{Reg}(T)$ $\leq 8^{3}K^{3}(1-p)^{-2}\log\frac{2KT^{2}}{\delta}+\sum_{t\in[T]\setminus\mathcal{E}_{T}}\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]$
$=$ $8^{3}K^{3}(1-p)^{-2}\log\frac{2KT^{2}}{\delta}+\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}\left(a_{t}=\widehat{a}_{t}\right)\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}$
@@ E.3 Proof of Theorem 3
$\leq\frac{16\sigma\widetilde{\sigma}_{\max}}{p\widetilde{\sigma}_{\min}}\sqrt{\frac{2(d+d_{h})}{t}\log\frac{2Kt^{2}}{\delta}},$
@@
$\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}\left(a_{t}=\widehat{a}_{t}\right)\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}\leq\frac{16\sigma\widetilde{\sigma}_{\max}}{p\widetilde{\sigma}_{\min}}\sqrt{2(d+d_{h})T\log\frac{2KT^{2}}{\delta}}.$
By Lemma 5 and the Hoeffding bound, with probability at least $1-\delta$
$\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}\left(a_{t}\neq\widehat{a}_{t}\right)\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}$
$\leq 2\left[\sum_{t\in[T]}\mathbb{I}\left(a_{t}\neq\widehat{a}_{t}\right)-\mathbb{P}\left(a_{t}\neq\widehat{a}_{t}\right)+\mathbb{P}\left(a_{t}\neq\widehat{a}_{t}\right)\right]$
$\leq 4\sqrt{2T\log\frac{2}{\delta}}+\frac{4p\sqrt{T}}{K-1}+4\delta.$
Then with probability at least $1-6\delta$,
$\mathrm{Reg}(T)$ $\leq 8^{3}K^{3}(1-p)^{-2}\log\frac{2KT^{2}}{\delta}+\frac{4\sqrt{T}}{K-1}\frac{\log((T+1)^{2}/\delta)}{\log(1/p)}$
$\quad+4\sqrt{2T\log\frac{2}{\delta}}+4\delta+\frac{16\sigma\widetilde{\sigma}_{\max}}{p\widetilde{\sigma}_{\min}}\sqrt{2(d+d_{h})T\log\frac{2KT^{2}}{\delta}},$
@@ E.4 Proof of Theorem 4
$\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star})=\widetilde{\mathbf{x}}_{a}^{\top}\widetilde{\mathbf{V}}_{t}^{-1}\left\{\sum_{\tau=1}^{t}\mathbb{I}(\mathcal{M}_{\tau})\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\left(\widetilde{y}_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}{\boldsymbol{\mu}}_{\star}\right)-{\boldsymbol{\mu}}_{\star}\right\}.$
@@
$\widetilde{y}_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}{\boldsymbol{\mu}}_{\star}=\left(1-\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,t}}\right)\widetilde{\mathbf{x}}_{a}^{\top}\left(\check{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star}\right)+\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\epsilon_{\tau}.$
@@
$\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star})=\widetilde{\mathbf{x}}_{a}^{\top}\widetilde{\mathbf{V}}_{t}^{-1}\left\{\left(\widetilde{\mathbf{V}}_{t}-\widetilde{\mathbf{A}}_{t}\right)\left(\check{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star}\right)+\sum_{\tau=1}^{t}\mathbb{I}(\mathcal{M}_{\tau})\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\widetilde{\mathbf{x}}_{a}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right\}.$
@@
$\check{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star}=$ $\left(\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}+p\mathbf{I}_{K}\right)^{-1}\left(\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{\tau}-p{\boldsymbol{\mu}}_{\star}\right)$
$=$ $\left(\sum_{\tau=1}^{t}\frac{1}{\phi_{a_{\tau},\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K}\right)^{-1}\left(\sum_{\tau=1}^{t}\frac{1}{\phi_{a_{\tau},\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right)$
$=\left(\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K}\right)^{-1}\left(\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right),$
@@
$\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star})=$ $\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1}\left\{\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{A}_{t}^{-1}\left(\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right)+\sum_{\tau=1}^{t}\frac{\epsilon_{\tau}}{\phi_{a_{\tau},\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}-{\boldsymbol{\mu}}_{\star}\right\}$
$=$ $\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1}\left\{\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{A}_{t}^{-1}+\mathbf{I}_{K}\right\}\left(\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right)$
$=$ $\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}^{1/2}\mathbf{A}_{t}^{-1}\mathbf{V}_{t}^{1/2}\right)\mathbf{V}_{t}^{-1/2}\left(\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right).$
@@
$\max_{a\in[K]}|{\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star})}|\leq\max_{a\in[K]}\|\widetilde{\mathbf{x}}_{a}\|_{\mathbf{V}_{t}^{-1}}\|\mathbf{V}_{t}^{1/2}\mathbf{A}_{t}^{-1}\mathbf{V}_{t}^{1/2}\|_{2}\left\|\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right\|_{\mathbf{V}_{t}^{-1}}.$
@@
$\max_{a\in[K]}|{\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star})}|$ $\leq\frac{\max_{a\in[K]}\|\widetilde{\mathbf{x}}_{a}\|_{\mathbf{V}_{t}^{-1}}}{1-\epsilon}\left\|\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right\|_{\mathbf{V}_{t}^{-1}}$
$\leq\frac{\max_{a\in[K]}\|\widetilde{\mathbf{x}}_{a}\|_{\mathbf{V}_{t}^{-1}}}{1-\epsilon}\left(\frac{1}{p}\left\|\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{\tau}\right\|_{\mathbf{V}_{t}^{-1}}+\left\|{\boldsymbol{\mu}}_{\star}\right\|_{\mathbf{V}_{t}^{-1}}\right).$
Note that the matrix $\mathbf{V}_{t}$ is deterministic. By Lemma 9 in (Abbasi-yadkori et al., 2011), with probability at least $1-\delta,$
$\left\|\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{\tau}\right\|_{\mathbf{V}_{t}^{-1}}$ $\leq\left\|\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{\tau}\right\|_{\left(\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K}\right)^{-1}}$
@@
$\left\|\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{\tau}\right\|_{\mathbf{V}_{t}^{-1}}\leq\sigma\sqrt{K\log\frac{t+1}{\delta}},$
@@
$\max_{a\in[K]}|{\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star})}|$ $\leq\frac{\max_{a\in[K]}\|\widetilde{\mathbf{x}}_{a}\|_{\mathbf{V}_{t}^{-1}}}{1-\epsilon}\left(\frac{\sigma}{p}\sqrt{K\log\frac{t+1}{\delta}}+\left\|{\boldsymbol{\mu}}_{\star}\right\|_{\mathbf{V}_{t}^{-1}}\right)$
@@
E.5 Proof of Theorem 5
Because the regret is bounded by $1$ and the number of rounds for the exploration phase is at most $|\mathcal{E}_{T}|\leq 32(1-p)^{-2}K^{2}\log({2dT^{2}}/{\delta}).$